# Differentially Private Graph Diffusion with Applications in Personalized PageRanks

### v1

@@ 4 Experiments
In this section, we present empirical evaluations to support our theoretical findings. Specifically, we apply the widely-used PPR algorithm (Sec. 3.4) to real-world graphs. We focus on the accuracy of noisy PPR in ranking tasks under personalized edge-level DP due to its practicality as noted in [20].
@@ 4.2 Ablation Study
Degree-based Thresholding Function & $\bm{\infty}$-Wasserstein Tracking. We evaluate the effectiveness of our graph-dependent thresholding function $f$ and $\infty$-Wasserstein distance tracking. We vary $\eta$ uniformly across seven discrete values in from $10^{-4}$ to $10^{-10}$ and report the optimal performance for all methods in Fig. 6. Firstly, we compare the degree-based thresholding function $f(\mathbf{s})=\min(\max(\mathbf{x},0),\eta\cdot\mathbf{d})$ (Red $\bullet$) against its graph-independent variant $\tilde{f}(\mathbf{s})=\min(\max(\mathbf{x},0),\eta\cdot\bm{1})$ (Red $\times$), which uniformly clips over nodes regardless of their degree. As $\epsilon$ varies from $0.1$ to $3$, the degree-based thresholding function consistently outperforms its graph-independent counterpart, with a margin of $0.1$ to $0.16$ in the NDCG score. This verifies the effectiveness of graph-dependent thresholding function $f$, which can balance the privacy-utility trade-off more effectively by focusing less on sensitive low-degree nodes. Note that the graph-dependent (in)dependent thresholding function $f$ ($\tilde{f}$) naturally induces a diameter $\eta|\mathcal{V}|$ ($2\eta|\mathcal{E}|$), respectively, which can be utilized in the modification of Theorem 1 for privacy accounting. However, the graph dependent thresholding function $f$ results in a significantly larger diameter, which substantially degrades utility. This underscores the importance of leveraging $\infty$-Wasserstein tracking.
@@ 5 Conclusion
Societal Impact and Limitations. Our work advances the development of DP graph algorithms, offering strong privacy protection when properly implemented. At present, we do not aware of any negative social impact resulting from this research. The effectiveness of our noisy graph diffusion framework has been demonstrated primarily in the context of PPR diffusion. Extending these results to other types of graph diffusions, such as heat kernel diffusion, is a promising direction for future research. This paper focuses on edge-level privacy protections; exploring extensions to node-level DP protections could further broaden the scope of our research.
@@ Appendix A Preliminaries
Appendix B Privacy Guarantees for General Noisy Graph Diffusion: Analytical Proofs
In this section, we first extend our noisy graph diffusion to accommodate general noise allocations and provide the corresponding privacy guarantees. We then present the main proof for these results. Additionally, we consider uniform noise allocation as a special case within this framework and derive the results for Theorem 1 as detailed in Sec. 3.
B.1 Noisy Graph Diffusion with General Noise Allocations
In Sec. 3, we mainly considers noisy graph diffusion framework based on uniform noise scheduling while in this section, we introduce noisy graph diffusions with general noise allocations. Specifically, we introduce our noisy graph diffusion framework as follows:
Given a graph diffusion $\mathscr{D}=\{\phi_{k}\}_{k=1}^{\infty}$ and a total diffusion step $K$, the general noisy graph diffusion $\mathscr{D}_{K,\Pi}$ is defined by a series of composed noisy graph diffusion mappings $\phi_{k,\sigma_{k}}$, noise-free diffusion mappings $\phi_{l,f}$, and a noise allocation tuple $\Pi=(m,\sigma_{m+1},\sigma_{m+2},\ldots,\sigma_{K})$:
$\mathscr{D}_{K,\Pi}=\phi_{K,\sigma_{K}}\circ\phi_{K-1,\sigma_{K-1}}\circ\cdots\circ\phi_{m+1,\sigma_{m+1}}\circ\phi_{m,f}\circ\cdots\circ\phi_{1,f}$ (16)
where $\phi_{k,\sigma_{k}}(\mathbf{s}_{k-1})=\phi_{k}(f(\mathbf{s}_{k-1}))+\xi_{k}^{(1)}+\xi_{k}^{(2)}$ and $\phi_{l,f}(\mathbf{s}_{l-1})=\phi_{l}(f(\mathbf{s}_{l-1}))$ with the degree-based function $f(\mathbf{x})=\min(\max(\mathbf{x},0),\eta\cdot d)$. The noise allocation tuple determines when to introduce randomness and the noise scales at each step.
The privacy guarantees of the noisy graph diffusion with general noise allocations, as defined in Eq. (16), are summarized in the following theorem:
Theorem B.1 (Privacy Guarantees of Noisy Graph Diffusions).
Given a graph $\mathcal{G}$, an associate graph diffusion $\mathscr{D}=\{\phi_{k}\}_{k=1}^{\infty}$ and a noise allocation tuple $\Pi=(m,\sigma_{m+1},\cdots,\sigma_{K})$, define degree-based thresholding function $f(\mathbf{x})=\min(\max(\mathbf{x},0),\eta\cdot d),\forall k\in\{m+2,...,K\}$, then noisy graph diffusion mechanism $\mathscr{D}_{K,\Pi}$ satisfies edge-level $(\alpha,\epsilon)$-RDP with $\epsilon$ satisfies:
$\epsilon=\min_{\tau\in\{m,...,K-1\}}\left[g_{\alpha}(\sigma_{K},\gamma_{\text{max}}^{K-\tau}w_{\tau})+\sum_{k=\tau+1}^{K}g_{\alpha}(\sigma_{k},\rho_{\text{diff}}^{\mathbbm{1}_{k\neq m+1}}\lambda_{m+1}^{1-\mathbbm{1}_{k\neq m+1}})\right]$ (17)
where $g_{\alpha}(\sigma,\rho)=\frac{1}{\alpha-1}\ln(\frac{\alpha}{2\alpha-1}\exp(\frac{\alpha-1}{\sigma}\rho)+\frac{\alpha-1}{2\alpha-1}\exp(-\frac{\alpha}{\sigma}\rho))$, the distortion $\rho_{\text{diff}}=(4\gamma_{\text{max}}^{(1)}+2\gamma_{\text{max}})\eta$, noise-free shift $\lambda_{m+1}=\frac{(1-\gamma_{\text{max}}^{m+1})\rho_{\text{diff}}}{(1-\gamma_{\text{max}})}$, and distortion $w_{\tau}=\frac{\rho_{\text{diff}}\cdot(1-\gamma_{\text{max}}^{\tau})}{1-\gamma_{\text{max}}}$ with Lipschitz constant $\gamma_{\text{max}}=\max_{k\in[K]}\sum_{j=1}^{2}|\gamma_{j,k}|$, and maximum diffusion coefficient $\gamma_{\text{max}}^{(1)}=\max_{k\in[K]}|\gamma_{1,k}|$.
B.2 Main Proof
Proof of Theorem B.1.
@@ B.2 Main Proof
Given a seed $\mathbf{s}$, a total diffusion step $K$, a noise allocation tuple $\Pi=(0,\sigma_{m+1},...,\sigma_{K})$, and thresholding functions $f,f^{\prime}$, we consider two coupled graph diffusions $\mathscr{D}_{K,\Pi}(\mathbf{s}),\mathscr{D}_{K,\Pi}^{\prime}(\mathbf{s})$ which propogate on two joint edge-level adjacent graphs $\mathcal{G},\mathcal{G}^{\prime}$ respectively with diffusion mapping $\phi_{k},\phi_{k}^{\prime},k\in[K]$. Specifically, we have
$\mathscr{D}_{K,\Pi}(\mathbf{s}):\quad$ $\mathbf{s}_{k}=\phi_{k}(f(\mathbf{s}_{k-1}))+\xi_{k}^{(1)}+\xi_{k}^{(2)},m+1\leq k\leq K,$
$\mathbf{s}_{l}=\phi_{l}(f(\mathbf{s}_{l-1})),1\leq l\leq m;$ (18)
$\mathscr{D}_{K,\Pi}^{\prime}(\mathbf{s}):\quad$ $\mathbf{s}_{k}^{\prime}=\phi_{k}^{\prime}(f^{\prime}(\mathbf{s}_{k-1}^{\prime}))+\xi_{k}^{\prime(1)}+\xi_{k}^{\prime(2)},m+1\leq k\leq K,$
$\mathbf{s}_{l}^{\prime}=\phi_{l}^{\prime}(f^{\prime}(\mathbf{s}_{l-1}^{\prime})),1\leq l\leq m..$ (19)
where $\xi_{k}^{(1)},\xi_{k}^{\prime(1)},\xi_{k}^{(2)},\xi_{k}^{\prime(2)}$ are distributed according to the law $\mathcal{L}(\mathbf{0},\sigma_{k})$.
@@
Step 1: Distortion of Single-Step Graph Diffusion In each diffusion step, a distortion over the diffusion mappings $\phi_{k}\circ f$ is introduced due to the graph adjacency between the graph diffusions $\mathscr{D}_{K,\Pi}(\mathbf{s})$ and $\mathscr{D}_{K,\Pi}^{\prime}(\mathbf{s})$. This distortion, which is crucial for the subsequent analysis, is characterized by the following lemma:
@@
Given two graph diffusions $\mathscr{D}_{K,\Pi}(\mathbf{s}),\mathscr{D}_{K,\Pi}^{\prime}(\mathbf{s})$ mentioned above. Let $f,f^{\prime}$ denote corresponding thresholding operators. The diffusion distortion:
$\sup_{\mathbf{x}\in\mathbb{R}^{n}}\|\phi_{k}(f(\mathbf{x}))-\phi_{k}^{\prime}(f^{\prime}(\mathbf{x}))\|_{1}\leq 4\gamma_{\text{max}}^{(1)}\cdot\max_{\mathbf{x}\in\mathbb{R}^{n}}\left|\frac{[f(\mathbf{x})]_{i}}{d_{i}}\right|+\gamma_{\text{max}}\cdot\Delta f$ (20)
where $\Delta f=\sup_{x\in\mathbb{R}^{n}}\|f(x)-f^{\prime}(x)\|_{1}$, $\gamma_{\text{max}}=\max_{k}\sum_{j=1}^{2}|\gamma_{j,k}|$, and $\gamma_{\text{max}}^{(1)}=\max_{k}|\gamma_{1,k}|$.
We define the distortion as $h=4\gamma_{1}\cdot\max_{\mathbf{x}\in\mathbb{R}^{n}}\left|\frac{[f(\mathbf{x})]_{i}}{d_{i}}\right|+\gamma_{\text{max}}\cdot\Delta f$. The $\ell_{1}$ sensitivity between the thresholding functions $f$ and $f^{\prime}$ over edge-adjacent graphs, i.e., $\Delta f$, is easily controlled since only one edge differs. Therefore, the dominant term to control in the distortion is $\max_{\mathbf{x}\in\mathbb{R}^{n},i}\left|\frac{[f(\mathbf{x})]_{i}}{d_{i}}\right|$. When selecting $f$ as the degree-based thresholding function, i.e., $f(\mathbf{x})=\min(\max(\mathbf{x},0),\eta\cdot\mathbf{d})$, the distortion $h$ can be further simplified to $\rho_{\text{diff}}=(4\gamma_{\text{max}}^{(1)}+2\gamma_{\text{max}})\eta$.
Step 2: Shift of Graph Diffusions without Randomness. Further, we examine the coupled diffusion processes. For diffusion step from $l\in[m]$, as shown in Eq. (18) and Eq. (19), no randomness is injected in the iterates. However, since $\mathcal{G}$ and $\mathcal{G}^{\prime}$ are edge-level adjacent, the difference in $\phi_{k}$ and $\phi_{k}^{\prime}$ lead to a distortion in each diffusion step. Therefore, at time step $m$, two processes are different from each other with a shift described by the following lemma:
Lemma B.2 (Shift of Graph Diffusions without Randomness).
Consider two graph diffusions $\mathscr{D}_{K,\Pi}(\mathbf{s})$ and $\mathscr{D}_{K,\Pi}^{\prime}(\mathbf{s})$ as previously described. The difference in their states at any step $m$ is bounded as follows:
$\|\mathbf{s}_{m}-\mathbf{s}_{m}^{\prime}\|_{1}\leq\frac{(1-\gamma_{\text{max}}^{m})\rho_{\text{diff}}}{(1-\gamma_{\text{max}})}.$ (21)
Step 3: Bounding the Privacy Loss via Shift Absorption and PABI. Further, we consider the noisy diffusion process, where we introduce Laplacian randomness during graph propagation. For $k\geq m+1$, we can regard these two diffusion processes as coupled CNIs starting from different initial points with maximum shift $\frac{(1-\gamma_{\text{max}}^{m})\rho_{\text{diff}}}{(1-\gamma_{\text{max}})}$. If both processes share identical diffusion mappings, then according to the PABI principle, privacy guarantees will be amplified at an exponential rate. With this in mind, and drawing upon the noise-splitting approach outlined in [27], we mitigate the shifts caused by diffusion and thresholding by integrating these factors into the noise distribution. Consequently, we construct conditional CNI sequences with identical diffusion mappings for both processes. Specifically, we reformulate the diffusion process $\mathscr{D}_{K,\Pi}^{\prime}(\mathbf{s})$ as follows:
@@
where we introduce $\tilde{\xi}_{k}^{\prime(2)}$ as shifted Laplacian noise. Therefore, the coupled graph diffusion processes can be summarized as follows:
$\mathbf{s}_{k}=\underbrace{\phi_{k}(f(\mathbf{s}_{k-1}))+\xi_{k}^{(1)}}_{\text{Identical CNI}}+\xi_{k}^{(2)},\;\mathbf{s}_{k}^{\prime}=\underbrace{\phi_{k}(f(\mathbf{s}_{k-1}^{\prime}))+\xi_{k}^{\prime(1)}}_{\text{Identical CNI}}+\tilde{\xi}_{k}^{\prime(2)},\;\forall k\geq m+1.$ (23)
Note that our objective is to establish an upper bound for $\mathcal{D}_{\alpha}(\mathbf{s}_{K}\|\mathbf{s}_{K}^{\prime})$, we claim that for any parameter $\tau\in\{m,m+1,...,K-1\}$,
@@
As demonstrated in Eq. 25, the privacy leakage can be upper bounded by the Rényi divergence across joint Laplacian distributions with a shift (shift absorption term) and the divergence across conditional CNIs employing identical transformations $\phi_{k}\circ f$ (PABI term). The parameter $\tau$ is introduced to balance the privacy leakage from shifts between noise distributions against the leakage from CNIs starting from different initial conditions. Both terms can be further bounded as detailed in the following lemmas. It is important to emphasize that, for the latter, we leverage $\infty$-Wasserstein distance tracking method that get rid of the diameter requirement in original PABI analysis.
@@
Given two coupled graph diffusions mentioned above, for any $\tau\geq m$, the shift absorption term can be upper bounded by
$\mathcal{D}_{\alpha}(\xi_{\tau+1:K}^{(2)}\|\tilde{\xi}_{\tau+1:K}^{\prime(2)})\leq\sum_{k=\tau+1}^{K}g_{\alpha}(\sigma_{k},\rho_{\text{diff}}^{\mathbbm{1}_{k\neq m+1}}\lambda_{m+1}^{1-\mathbbm{1}_{k\neq m+1}})$ (26)
where distortion $\rho_{\text{diff}}=(4\gamma_{\text{max}}^{(1)}+2\gamma_{\text{max}})\eta$, noise-free shift $\lambda_{m+1}=\frac{(1-\gamma_{\text{max}}^{m+1})\rho_{\text{diff}}}{(1-\gamma_{\text{max}})}$ and shifted laplace function $g_{\alpha}(\sigma,\rho)=\frac{1}{\alpha-1}\ln(\frac{\alpha}{2\alpha-1}\exp(\frac{\alpha-1}{\sigma}\rho)+\frac{\alpha-1}{2\alpha-1}\exp(-\frac{\alpha}{\sigma}\rho))$.
@@
Given two coupled graph diffusions mentioned above, for any $\tau\geq m$, we have
@@
Given two coupled graph diffusions mentioned above, for any $\tau\geq m$, we have
@@
where distortion $\rho_{\text{diff}}=(4\gamma_{\text{max}}^{(1)}+2\gamma_{\text{max}})\eta$.
By summarizing the above results (Lemma B.3B.4, and B.5), we conclude the final results:
@@
$\leq$ $\sum_{k=\tau+1}^{K}g_{\alpha}(\sigma_{k},\rho_{\text{diff}}^{\mathbbm{1}_{k\neq m+1}}\lambda_{m+1}^{1-\mathbbm{1}_{k\neq m+1}})+g_{\alpha}(\sigma_{K},\gamma_{\text{max}}^{K-\tau}w_{\tau})$ (29)
B.3 Proof of Theorem 1
According to the proof in Sec. B.2, by setting $m=0$, we have $\lambda_{1}=\rho_{\text{diff}}$. Furthermore, since $w_{\tau}\leq\frac{\rho_{\text{diff}}}{1-\gamma_{\text{max}}}$, by selecting a uniform noise allocation schedule, i.e., $\sigma_{k}=\sigma$, we obtain the result in Theorem 1:
$\epsilon=\min_{\tau\in\{0,...,K-1\}}\left[g_{\alpha}(\sigma,\frac{\rho_{\text{diff}}\cdot\gamma_{\text{max}}^{K-\tau}}{1-\gamma_{\text{max}}})+\sum_{k=\tau+1}^{K}g_{\alpha}(\sigma,\rho_{\text{diff}})\right]$ (30)

### v4

@@ 4 Experiments
In this section, we present empirical evaluations to support our theoretical findings. Specifically, we apply the widely-used PPR algorithm (Sec. 3.4) to real-world graphs. In practice, we also include a projection step onto the unit $\ell_{1}$ ball after injecting Laplace noise at each diffusion step. This adjustment has been observed to slightly improve the utility of the resulting PPR without impacting our theoretical analysis (see App. B.1 for details). We focus on the accuracy of noisy PPR in ranking tasks under personalized edge-level DP due to its practicality as noted in [20].
@@ 4.2 Ablation Study
Degree-based Thresholding Function & $\bm{\infty}$-Wasserstein Tracking. We evaluate the effectiveness of our graph-dependent thresholding function $f$ and $\infty$-Wasserstein distance tracking. We vary $\eta$ uniformly across seven discrete values in from $10^{-4}$ to $10^{-10}$ and report the optimal performance for all methods in Fig. 6. Firstly, we compare the degree-based thresholding function $f(\mathbf{s})=\min(\max(\mathbf{x},0),\eta\cdot\mathbf{d})$ (Red $\bullet$) against its graph-independent variant $\tilde{f}(\mathbf{s})=\min(\max(\mathbf{x},0),\eta\cdot\bm{1})$ (Red $\times$), which uniformly clips over nodes regardless of their degree. As $\epsilon$ varies from $0.1$ to $3$, the degree-based thresholding function consistently outperforms its graph-independent counterpart, with a margin of $0.15$ to $0.2$ in the NDCG score. This verifies the effectiveness of graph-dependent thresholding function $f$, which can balance the privacy-utility trade-off more effectively by focusing less on sensitive low-degree nodes.
@@
Note that the graph-dependent (in)dependent thresholding function $f$ ($\tilde{f}$) naturally induces a diameter $\eta|\mathcal{V}|$ ($2\eta|\mathcal{E}|$), respectively, which can be utilized in the modification of Theorem 1 for privacy accounting. However, the graph dependent thresholding function $f$ results in a significantly larger diameter, which substantially degrades utility. This underscores the importance of leveraging $\infty$-Wasserstein tracking.
@@ 5 Conclusion
Societal Impact and Limitations. Our work advances the development of DP graph algorithms, offering strong privacy protection when properly implemented. For its limitations, we refer readers to standard textbooks on the subject [42]. The effectiveness of our noisy graph diffusion framework has been demonstrated primarily in the context of PPR diffusion. Extending these results to other types of graph diffusions, such as heat kernel diffusion, is a promising direction for future research. This paper focuses on edge-level privacy protections; exploring extensions to node-level DP protections could further broaden the scope of our research.
@@ Appendix A Preliminaries
Appendix B Privacy Guarantees for Noisy Graph Diffusion: Analytical Proofs
In this section, we present the main proof for the results in Theorem 1 as detailed in Sec. 3.
B.1 Main Proof
Proof of Theorem 1.
@@ B.1 Main Proof
Given a seed $\mathbf{s}$, a total diffusion step $K$, a noise scale $\sigma$, and degree-based thresholding functions $f,f^{\prime}$, we consider two coupled graph diffusions $\mathscr{D}_{K,\sigma}(\mathbf{s}),\mathscr{D}_{K,\sigma}^{\prime}(\mathbf{s})$ which propogate on two joint edge-level adjacent graphs $\mathcal{G},\mathcal{G}^{\prime}$ respectively with diffusion mapping $\phi_{k},\phi_{k}^{\prime},k\in[K]$. Specifically, we have
$\mathscr{D}_{K,\sigma}(\mathbf{s}):\quad$ $\mathbf{s}_{k}=\phi_{k}(f(\mathbf{s}_{k-1}))+\xi_{k}^{(1)}+\xi_{k}^{(2)},1\leq k\leq K,$ (16)
$\mathscr{D}_{K,\sigma}^{\prime}(\mathbf{s}):\quad$ $\mathbf{s}_{k}^{\prime}=\phi_{k}^{\prime}(f^{\prime}(\mathbf{s}_{k-1}^{\prime}))+\xi_{k}^{\prime(1)}+\xi_{k}^{\prime(2)},1\leq k\leq K.$ (17)
where $\xi_{k}^{(1)},\xi_{k}^{\prime(1)},\xi_{k}^{(2)},\xi_{k}^{\prime(2)}$ are distributed according to the law $\mathcal{L}(\mathbf{0},\sigma)$.
@@
Step 1: Distortion of Single-Step Graph Diffusion In each diffusion step, a distortion over the diffusion mappings $\phi_{k}\circ f$ is introduced due to the graph adjacency between the graph diffusions $\mathscr{D}_{K,\sigma}(\mathbf{s})$ and $\mathscr{D}_{K,\sigma}^{\prime}(\mathbf{s})$. This distortion, which is crucial for the subsequent analysis, is characterized by the following lemma:
@@
Given two graph diffusions $\mathscr{D}_{K,\sigma}(\mathbf{s}),\mathscr{D}_{K,\sigma}^{\prime}(\mathbf{s})$ mentioned above. Let $f,f^{\prime}$ denote corresponding graph-dependent degree-based thresholding operators, i.e., $f(\mathbf{x})=\min(\max(\mathbf{x},-\eta\cdot\mathbf{d}),\eta\cdot\mathbf{d})$. The diffusion distortion satisfies:
$\sup_{\mathbf{x}\in\mathbb{R}^{n}}\|\phi_{k}(f(\mathbf{x}))-\phi_{k}^{\prime}(f^{\prime}(\mathbf{x}))\|_{1}\leq\max(4\gamma_{\text{max}}^{(1)},2\gamma_{\text{max}})\cdot\eta$ (19)
where Lipschitz constant $\gamma_{\text{max}}=\max_{k}|\gamma_{1,k}|+|\gamma_{2,k}|$, and maximum diffusion coefficient $\gamma_{\text{max}}^{(1)}=\max_{k}|\gamma_{1,k}|$.
We denote this single-step graph diffusion distortion bound as $\rho_{\text{diff}}$, i.e., $\rho_{\text{diff}}=\max(4\gamma_{\text{max}}^{(1)},2\gamma_{\text{max}})\cdot\eta$. Moreover, incorporating an $\ell_{1}$-ball projection $\mathscr{P}_{\mathcal{B}}$ into the pipeline, i.e., adopting $\phi_{k}\circ f\circ\mathscr{P}_{\mathcal{B}}$, does not alter this distortion bound.
According to the proof, this distortion analysis is tight and conveys several key insights. First, when the graph diffusion process diffuses relatively slow, i.e., $\gamma_{1,k}<\gamma_{2,k}$, the distortion is tight and primarily governed by the Lipschitz constant of the diffusion mapping. In contrast, when the diffusion is relatively fast, the distortion is controlled by the maximum diffusion coefficient $\gamma_{1,k}$. In this latter scenario, the bound becomes asymptotically tight when the graph structure satisfies the condition that the nodes connected by the perturbed edge have no common neighbors, and their degrees tend to infinity. We raise a double star graph as an toy example that satisfies the above no common neighbor condition (see Fig. 7).
Figure 7: Double Star Graph.
Step 3: Bounding the Privacy Loss via Shift Absorption and PABI. Next, we consider the noisy diffusion process, where Laplacian noise is introduced during graph propagation. Note that, we focus on injecting noise at the initial step of the diffusion process, although the analysis can be directly extended to include noise injection at intermediate steps. For $k\geq 1$, drawing upon the noise-splitting approach outlined in [27], we mitigate the shifts caused by diffusion and thresholding by integrating these factors into the noise distribution. Consequently, we construct conditional CNI sequences with identical diffusion mappings for both processes. Specifically, we reformulate the diffusion process $\mathscr{D}_{K,\Pi}^{\prime}(\mathbf{s})$ as follows:
@@
where we introduce $\tilde{\xi}_{k}^{\prime(2)}$ as shifted Laplacian noise and noise scale $\sigma_{k}=\sigma$. Therefore, the coupled graph diffusion processes can be summarized as follows:
$\mathbf{s}_{k}=\underbrace{\phi_{k}(f(\mathbf{s}_{k-1}))+\xi_{k}^{(1)}}_{\text{Identical CNI}}+\xi_{k}^{(2)},\;\mathbf{s}_{k}^{\prime}=\underbrace{\phi_{k}(f(\mathbf{s}_{k-1}^{\prime}))+\xi_{k}^{\prime(1)}}_{\text{Identical CNI}}+\tilde{\xi}_{k}^{\prime(2)},\;\forall k\geq 1.$ (21)
Note that our objective is to establish an upper bound for $\mathcal{D}_{\alpha}(\mathbf{s}_{K}\|\mathbf{s}_{K}^{\prime})$, we claim that for any parameter $\tau\in\{0,1,...,K-1\}$,
@@
As demonstrated in Eq. 23, the privacy leakage can be upper bounded by the Rényi divergence across joint Laplacian distributions with a shift (shift absorption term) and the divergence across conditional CNIs employing identical transformations $\phi_{k}\circ f$ (PABI term). The parameter $\tau$ is introduced to balance the privacy leakage from shifts between noise distributions against the leakage from CNIs starting from different initial conditions. Both terms can be further bounded as detailed in the following lemmas. It is important to emphasize that, for the latter, we leverage $\infty$-Wasserstein distance tracking method that get rid of the diameter requirement in original PABI analysis. Besides, for each lemma below, we state that considering projection onto the $\ell_{1}$ ball does not affect the conclusions.
@@
Given two coupled graph diffusions mentioned above, for any $\tau\geq 0$, the shift absorption term can be upper bounded by
$\mathcal{D}_{\alpha}(\xi_{\tau+1:K}^{(2)}\|\tilde{\xi}_{\tau+1:K}^{\prime(2)})\leq\sum_{k=\tau+1}^{K}g_{\alpha}(\sigma_{k},\rho_{\text{diff}})$ (24)
where distortion $\rho_{\text{diff}}=\max(4\gamma_{\text{max}}^{(1)},2\gamma_{\text{max}})\cdot\eta$, and shifted Laplace function $g_{\alpha}(\sigma,\rho)=\frac{1}{\alpha-1}\ln(\frac{\alpha}{2\alpha-1}\exp(\frac{\alpha-1}{\sigma}\rho)+\frac{\alpha-1}{2\alpha-1}\exp(-\frac{\alpha}{\sigma}\rho))$.
@@
Given two coupled graph diffusions mentioned above, for any $\tau\geq 0$, we have
@@
Given two coupled graph diffusions mentioned above, for any $\tau\geq 0$, we have
@@
where distortion $\rho_{\text{diff}}=\max(4\gamma_{\text{max}}^{(1)},2\gamma_{\text{max}})\cdot\eta$.
By summarizing the above results (Lemma B.2B.3, and B.4), we conclude the final results:
@@
$\leq$ $\sum_{k=\tau+1}^{K}g_{\alpha}(\sigma_{k},\rho_{\text{diff}})+g_{\alpha}(\sigma_{K},\gamma_{\text{max}}^{K-\tau}w_{\tau})$ (27)
$=$ $(K-\tau)\cdot g_{\alpha}(\sigma,\rho_{\text{diff}})+g_{\alpha}(\sigma,\gamma_{\text{max}}^{K-\tau}\cdot\frac{\rho_{\text{diff}}\cdot(1-\gamma_{\text{max}}^{\tau})}{1-\gamma_{\text{max}}})$ (28)